OpenStudy (anonymous):

Divide using synthetic division, and write a summary statement in fraction form. 2x^3 +3x^2 +4x - 10/ x+1

OpenStudy (jdoe0001):

do you know how to do synthetic division?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

idk sorry

OpenStudy (jdoe0001):

ok.... you'd grab the coefficients of the dividend and put them on the right-hand-side then the divisor will go on the left-hand side notice that the divisor is x+1, thus x+1 =0 => x= -1 so we'll end up using for our divisor, " -1 " only so you'd drop the 1st coefficient and multiply the divisor by it, the product goes underneath the next coefficient

OpenStudy (jdoe0001):

\(2x^3 +3x^2 +4x - 10 \div x+1\\ \quad \\ \begin{array}{rrrrrr} -1&|&2&3&4&-10\\ &|&&-2\\ \hline\\ &&2&1 \end{array}\) notice we drop the 1st coefficient, that is "2", -1 * 2 = -2, -2 goes underneath the next coefficient, 3

OpenStudy (jdoe0001):

then we sum them up, 3 - 2 = 1 as you can see above then we again, do the divisor multiplication of the result of the sum so -1 * 1 = -1 so \(2x^3 +3x^2 +4x - 10 \div x+1\\ \quad \\ \begin{array}{rrrrrr} -1&|&2&3&4&-10\\ &|&&-2&-1\\ \hline\\ &&2&1&3 \end{array}\)

OpenStudy (anonymous):

okay

OpenStudy (jdoe0001):

4 - 1 = 3 again, we multiply the divisor times the sum result, that is -1 * 3 and that will go under the next coefficient.... so what do you get?

OpenStudy (anonymous):

-7

OpenStudy (jdoe0001):

anyhow, lemme post it -1 * 3 = -3 so \(\begin{array}{rrrrrr} -1&|&2&3&4&-10\\ &|&&-2&-1&-3\\ \hline\\ &&2&1&3&-13 \end{array}\)

OpenStudy (jdoe0001):

well, notice is -10, not +10 so, -10 -3 = -13

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

so -1 doesn't work ? or is the numbers at the bottom the answer ?

OpenStudy (jdoe0001):

one thing to recall is that the dividend is firstly arranged in descending order, the 3rd power to the far-left and then down to the right, 2nd power, 1st power and then constant

OpenStudy (jdoe0001):

\(\bf 2x^\color{red}{3} +3x^\color{red}{2} +4x^\color{red}{1} - 10 \div x+1\)

OpenStudy (jdoe0001):

that's how the dividend should be arranged for any division really, either long or synthetic

OpenStudy (anonymous):

2x2 + x + 3 + (-13/x+1) 2x2 + 5x + 9 + (1/x+1) 2x2 + x - 3 + (13/x+1) 2x2 + 5x + 9 + (-1/x+1)

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

so from \(2x^3 +3x^2 +4x - 10 \div x+1\\ \quad \\ \begin{array}{rrrrrr} -1&|&2&3&4&-10\\ &|&&-2&-1&-3\\ \hline\\ &&2&1&3&-13 \end{array}\) we get a quotient, the quotient is those coefficients, with a lesser exponent than the original and the last number is our remainder

OpenStudy (anonymous):

so it would be a?

OpenStudy (jdoe0001):

\({\bf 2x^\color{red}{3} +3x^\color{red}{2} +4x - 10 \div x+1}\\ \quad \\ \begin{array}{rrrrrr} -1&|&2&3&4&-10\\ &|&&-2&-1&-3\\ \hline\\ &&2&1&3&-13 \end{array}\bf \implies 2x^\color{red}{2}+1x+3+\textit{(remainder here)}\) notice that our quotient function, has 1 degree less than the original, the original had a degree of 3, our quotient from the division will have 1 less, or 3-1 = 2, so starts off with an exponent of 2